How do you write 4.19 × 104 in standard form?

Grade:6

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given number is 4.19 X104
By multiplying 4.19 with 104we get the standard form
It can be written in standard form as 41900
So the standard form of the given expression can be written as 41,900
kartheeek
9 years ago
4.19*10^4=4.19*10000=41900